OpenStudy (anonymous):

2x-y=3 to get the ( y )by it self

OpenStudy (apoorvk):

add y-3 to both sides, and simplify!

OpenStudy (anonymous):

\[2x-y=3\] \[2x=y+3\] \[2x-3=y\]

OpenStudy (anonymous):

Thanks : )